A landscaping company places an order with a nursery for delivery of a total of 24 bushes.
Found 2 solutions by algebrahouse.com, richwmiller:
Answer by algebrahouse.com(1659)

s = number of small bushes
m = number of medium bushes

s + m = 24 {a total of 24 bushes were delivered}
23s + 47m = 864 {price of bush times number of bushes equals total price}

m = -s + 24 {subtracted s from each side of first equation}

23s + 47m = 864 {second equation}
23s + 47(-s + 24) = 864 {substituted (-s + 24), in for m, into second equation}
23s - 47s + 1128 = 864 {used distributive property}
-24s = -264 {subtracted 1128 from each side}
s = 11 {divided each side by -24}

m = -s + 24 {created equation above}
m = -11 + 24 {substituted 11, in for s, into created equation}
m = 13 {added}

11 small bushes were ordered
13 medium bushes were ordered

For more help like this, visit: www.algebrahouse.com



Answer by richwmiller(17219)

Linear_Equations_And_Systems_Word_Problems/1103316